# Can Large Language Models Improve Phishing Defense? A Large-Scale Controlled Experiment on Warning Dialogue Explanations

### v1

@@ Abstract
Phishing has become a prominent risk in modern cybersecurity, often used to bypass technological defences by exploiting predictable human behaviour. Warning dialogues are a standard mitigation measure, but the lack of explanatory clarity and static content limits their effectiveness. In this paper, we report on our research to assess the capacity of Large Language Models (LLMs) to generate clear, concise, and scalable explanations for phishing warnings. We carried out a large-scale between-subjects user study (N = 750) to compare the influence of warning dialogues supplemented with manually generated explanations against those generated by two LLMs, Claude 3.5 Sonnet and Llama 3.3 70B. We investigated two explanatory styles (feature-based and counterfactual) for their effects on behavioural metrics (click-through rate) and perceptual outcomes (e.g., trust, risk, clarity). The results indicate that well-constructed LLM-generated explanations can equal or surpass manually crafted explanations in reducing susceptibility to phishing; Claude-generated warnings exhibited particularly robust performance. Feature-based explanations were more effective for genuine phishing attempts, whereas counterfactual explanations diminished false-positive rates. Other variables such as workload, gender, and prior familiarity with warning dialogues significantly moderated warning effectiveness. These results indicate that LLMs can be used to automatically build explanations for warning users against phishing, and that such solutions are scalable, adaptive, and consistent with human-centred values.
@@
Phishing , Warnings , Human factors , Alert , LLMs , Explanations
journal: Computers & Security
orcid]https://orcid.org/$0$-0002-8261-3200
orcid]https://orcid.org/$0$-0001-9894-2116
orcid]https://orcid.org/$0$-0003-2730-7697
orcid]https://orcid.org/$0$-0001-7106-0463
orcid]https://orcid.org/$0$-0001-9916-271X
[…]
[unica]organization=University of Cagliari, addressline=Via Ospedale 72, city=Cagliari, postcode=09124, country=Italy
[…]
[uniba]organization=Department of Computer Science, University of Bari Aldo Moro, addressline=Via E. Orabona 4, city=Bari, postcode=70125, country=Italy
[…]
[kings]organization=King’s College London, addressline=Bush House, Strand Campus, 30, Aldwych, city=London, postcode=WC2b 4BG, country=UK
@@ 1 Introduction
Phishing attacks are among the most widespread and harmful cyber threats, leveraging human behaviour instead of technical weaknesses (Khonji et al., 2013; Prakash et al., 2010). There is a vast number of automatic filtering techniques (e.g., blacklists, machine learning-based detection). Still, these automatic defence systems are pretty far from perfect (Basit et al., 2021), and there is the strong possibility of false positives and, even more importantly, false negatives, which entail that users might receive harmful material.
@@
To overcome this issue, most defensive systems engage the human in the decision-making process by presenting users with warning dialogues, allowing them to make the final decision and possibly prevent a phishing attack (Desolda et al., 2019). Although warning dialogues are commonly used, they work well only when describing the risk and influencing the decision to act upon (IBM, 2024). Most dialogues used today are passive and mostly static, with no clear explanations as to why a particular message is harmful (Egelman et al., 2008; Wu et al., 2006).
@@
In the last few years, researchers working in Human-Centred Security and Explainable AI (XAI) have investigated whether feature-based explanations might be effective in mitigating phishing attacks (Guidotti et al., 2019b; Gunning and Aha, 2019). The clear advantage of such explanations is the reduction of susceptibility to phishing, but they are hand-made, expensive to maintain, and might not easily adjust to the changing circumstances of the threats (Desolda et al., 2023; Buono et al., 2023; Greco et al., 2025). Other ways of generating explanations have also been considered in general, such as the use of counterfactuals (Guidotti et al., 2019a) or the use of Large Language Models (LLMs), but they have not been investigated extensively in the specific field of cybersecurity (Xu et al., 2024).
@@
To bridge these gaps, we have carried out a large-scale, between-subjects user study ($N=750$). We investigated how well phishing warning dialogues with explanations can be generated with LLMs. We assessed the result of using a commercial closed-weight model (Claude 3.5 Sonnet) and an open-weight alternative (Llama 3.3 70B). In addition, we considered two styles of explanation, i.e., feature-based and counterfactual explanations. A baseline with feature-based, manually-written explanations has also been considered  (Desolda et al., 2023; Greco et al., 2025). The effectiveness of the the different experimental conditions have been evaluated by considering both measures of behaviour (e.g., click-through rate on phishing links) and user perceptions (e.g., clarity, trust, perceived risk), and modelled the degree to which individual factors, including cognitive style and knowledge of cybersecurity issues, influence results.
@@
We found that explanations generated by LLMs can be as effective as those produced manually; specifically, explanations generated by Claude were better than those crafted manually. Concerning the type of explanation, counterfactual ones have an interesting power in ensuring false positive cases, and feature-based ones have certain benefits in true positive cases. Moreover, user-related factors such as gender, familiarity, and mental workload significantly influence the effectiveness of the warning.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on phishing defences, warning dialogue design, and explainable AI in cybersecurity. Section 3 presents the experimental design and the explanation generation process. This section also reports the formalisation of the four research questions this study aims to answer. Section 4 details the study procedure and methodology. Section 5 reports the quantitative and qualitative results addressing our four research questions. Section 6 discusses threats to validity. Section 7 synthesises the findings as a set of lessons learned. In conclusion, Section  8 summarises the paper and emphasises potential areas for future research.
@@ 2.1 Warning dialogues for Phishing Defence
The complete elimination of phishing through automated filtering methods such as blocklists (Gupta and Kumaraguru, 2014) and AI-based methods Basit et al. (2021) is currently unfeasible. No existing tools can detect phishing content, such as websites or emails, with perfect accuracy (El Aassal et al., 2020; Mehdi Gholampour and Verma, 2023). Furthermore, attempts at reducing the number of false negatives (i.e., phishing emails that go undetected) tend to increase the number of false positives (i.e., legitimate emails misclassified as phishing) in classification tasks (Saxena, 2018). This eventually leads to filtering out genuine, relevant emails, potentially disrupting user productivity.
@@
The solution currently adopted in most of the systems is not to automatically filter suspicious content classified as phishing with lower confidence; instead, these emails are presented to users alongside a warning dialogue that alerts them about possible threats while leaving the final decision to the users (Kumaraguru et al., 2010). Despite warning dialogues, many users fail to grasp their significance or meaning, so phishing attacks remain highly effective (IBM, 2024).
@@
One key problem of everyday warning dialogues is their passive behaviour. Popular email clients such as Google Gmail implement static toolbars above the email content. Warnings like these are largely ineffective, as users often overlook or ignore them (Egelman et al., 2008; Wu et al., 2006). On the other hand, warnings that implement an active behaviour, i.e., interrupt user interactions and demand attention, have been proven to be significantly more effective (Wogalter et al., 2002; Petelka et al., 2019; Buono et al., 2023; Egelman et al., 2008; Greco et al., 2025). However, such warnings are currently implemented only in web browsers (Desolda et al., 2019), despite the benefits of active behaviour having also been demonstrated for warnings shown in email clients (Greco et al., 2025; Buono et al., 2023).
@@
Another limitation of warnings is the phenomenon known as habituation. This behaviour occurs when users are repeatedly exposed to the same warning stimulus, even under varying risk conditions, leading to decreased attentional responses and increased likelihood of bypassing the warning (Akhawe and Felt, 2013; Kim and Wogalter, 2009). To mitigate this problem, polymorphic warnings were demonstrated to be significantly more resistant to habituation than static warnings (Anderson et al., 2015). Such warnings are called polymorphic since they dynamically change their appearance or content to provide varied visual stimuli. Nonetheless, they are still not implemented in browsers or email clients despite their advantages (Desolda et al., 2019). Only two studies have explored using polymorphic warnings paired with tailored explanations to enhance their effectiveness (Desolda et al., 2023; Buono et al., 2023). Desolda et al. found that including manually designed explanations of malicious features in warning dialogues in web browsers led to warnings that were more familiar and understandable than those in the most common browsers. Buono et al. studied warnings with explanations in the context of email clients and found that active warnings with explanations based on phishing features largely outperformed contextual warnings without explanations. This study confirmed the importance of interrupting user interaction flow to focus entirely on the message, similar to active warnings shown in browsers. Additionally, the study demonstrated the importance of explaining to users why content is malicious, helping them make better-informed decisions.
@@
A further issue that compromises the warnings’ effectiveness is the absence of detailed explanations about the phishing risk (Bravo-Lillo et al., 2011). Ideally, warnings should communicate clearly why specific content is dangerous and outline the potential consequences of ignoring it. However, many current warnings provide vague messages that fail to identify the malicious elements within an email or website. This places the burden of risk assessment on the users, who often lack cybersecurity expertise or additional information to make informed decisions (Desolda et al., 2021). By incorporating explanations into warnings, the users’ understanding of the threat can be improved, addressing knowledge gaps, and potentially increasing compliance and trust in the system (Bravo-Lillo et al., 2011; Vilone and Longo, 2021).
@@
The use of explanatory warnings in web browsers has been explored in (Desolda et al., 2023). The authors identified seven website features that can be intuitively explained to users and produced two explanation messages for each. These messages were refined through an iterative process to optimise text readability and sentiment. Moreover, the explanation messages followed a well-defined structure grounded in warning theory (Bauer et al., 2013), which led to the identification of a template for generating explanation messages: “Feature description + Hazard Explanation + Consequences of not complying with the warning”. The authors empirically validated with 150 users the adoption of warnings with explanations, compared with warnings without explanations. Results showed that the proposed warnings were clearer and more effective than those implemented in modern browsers, which lack explanations. Greco et al. (2025) explored a similar approach in the context of email clients, combining active warning dialogues with embedded explanations. The findings indicated that this combination of active behaviour and explanation significantly outperformed the state-of-the-art email client warnings proposed by Petelka et al. (2019).
@@
As security software engineers, we must understand the underlying decision-making algorithm to generate explanations. In the case of AI-based phishing filtering methods, explanations can be produced by employing eXplainable AI (XAI) techniques (Guidotti et al., 2019b; Gunning and Aha, 2019). [^1] See (Viganò and Magazzeni, 2020) for a discussion on the relationship between explainable cybersecurity and explainable AI. These allow one to explain the behaviour and decisions of AI models, which are often black-box, and thus not directly interpretable even by AI engineers. With XAI techniques, for example, the system can provide a rationale for the reasons that led to the automatic classification of an email as phishing. Explanations in warning dialogues can contain such a rationale to improve the transparency of the AI system. Let us explore these aspects in the following subsection.
@@ 2.2 Human-Centred Explainable AI for Phishing Attacks
Integrating XAI solutions into cybersecurity systems is essential to maintain user trust and confidence. This underscores the importance of user-centred experimental evaluations that balance security, usability, and resilience to adversarial attacks (Charmet et al., 2022; Rjoub et al., 2023; Sarker et al., 2024; AL and Sagiroglu, 2025). In this regard, the broader Human-Centred XAI (HCXAI) literature offers valuable insights into multiple explanation types and their effects on users. The most prevalent explanations in AI-assisted decision systems are the feature-based ones (Lai et al., 2023), clarifying why an AI system produced a specific decision by highlighting the most important features driving the decision. Feature-based explanations might enhance user understanding, uncertainty awareness, trust calibration, and decision accuracy, though they can sometimes foster overreliance on AI (Zhang et al., 2020; Wang and Yin, 2021; Chen et al., 2023; Ma et al., 2023; Fok and Weld, 2024; Cau and Spano, 2025). Counterfactual explanations constitute another, less explored style: they provide not only why an AI system reached a specific decision, but also how it is possible for users to alter input features to achieve a potential desired outcome by presenting contrastive “what-if” statements. While they are mainly used to offer actionable recourse when negative decisions occur—e.g., loan denial—by indicating necessary feature changes for a positive result—e.g., increasing income for approval—(Koh et al., 2024; Verma et al., 2024; VanNostrand et al., 2024; Upadhyay et al., 2025), they have also demonstrated improvements in helpfulness, decision alignment, and accuracy, sometimes matching or exceeding feature-based methods, in AI-assisted decision-making tasks (Scharowski et al., 2023; Celar and Byrne, 2023; Cau et al., 2023; Teso et al., 2023; Lee and Chew, 2023; Gentile et al., 2025).
@@
Regarding XAI in cybersecurity and phishing in particular, only a small number of proof-of-concept studies have assessed explanation effectiveness with real users (Desolda et al., 2023; Greco et al., 2025). For example, Desolda et al. (2023) found that adding a feature-based explanation to a phishing warning dialogue made the warning more understandable and familiar, which may deter users from visiting malicious sites. In a similar study, Greco et al. (2025) compared warnings with and without feature-based explanations and the effect of showing the warning before versus after opening an email. They showed that adding explanations improved users’ comprehension of the phishing threat and, when the warning appeared after an email opening, led to a measurable reduction in click-through rate (CTR) despite the additional effort required by participants. However, these investigations have focused exclusively on manually crafted feature‐based explanations, leaving the generation of feature‐based or alternative explanation types via large language models largely unexplored (Sarker et al., 2024).

### v2

@@ Abstract
Phishing has become a prominent risk in modern cybersecurity, often used to bypass technological defences by exploiting predictable human behaviour. Warning dialogues are a standard mitigation measure, but the lack of explanatory clarity and static content limits their effectiveness. In this paper, we report on our research to assess the capacity of Large Language Models (LLMs) to generate clear, concise, and scalable explanations for phishing warnings. We carried out a large-scale between-subjects user study (N = 750) to compare the influence of warning dialogues supplemented with manually generated explanations against those generated by two LLMs, Claude 3.5 Sonnet and Llama 3.3 70B. We investigated two explanatory styles (feature-based and counterfactual) for their effects on behavioural metrics (click-through rate) and perceptual outcomes (e.g., trust, risk, clarity). The results provide empirical evidence that LLM-generated explanations achieve a level of protection statistically comparable to expert-crafted messages, effectively automating a high-cost task. While Claude 3.5 Sonnet showed a trend towards reducing click-through rates compared to manual baselines, Llama 3.3, despite being perceived as clearer, did not yield the same behavioral benefits. Feature-based explanations were more effective for genuine phishing attempts, whereas counterfactual explanations diminished false-positive rates. Other variables, such as workload, gender, and prior familiarity with warning dialogues, significantly moderated the effectiveness of warnings. These results indicate that LLMs can be used to automatically build explanations for warning users against phishing, and that such solutions are scalable, adaptive, and consistent with human-centred values.
@@
Phishing; Warnings; Human factors; Alert; LLMs; Explanations
@@ 1 Introduction
Phishing attacks are among the most widespread and harmful cyber threats, leveraging human behaviour instead of technical weaknesses (Khonji2013Phishing; Prakash2010PhishNet). There is a vast number of automatic filtering techniques (e.g., blacklists, machine learning-based detection), but these automatic defence systems are far from perfect (Basit2021Survey), and there is a strong possibility of false positives and, even more importantly, false negatives, which could result in users receiving harmful material.
@@
To overcome this issue, most defensive systems engage the human in the decision-making process by presenting users with warning dialogues, allowing them to make the final decision and possibly prevent a phishing attack (desolda2019alerting). Although warning dialogues are commonly used, they are effective only when clearly describing the risk and influencing the decision to act upon it (ibm2024security). Most dialogues used today are passive and mostly static, with no clear explanations as to why a particular message is harmful (egelman2008warned; wu2006security).
@@
In the last few years, researchers working in Human-Centred Security and Explainable AI (XAI) have investigated whether feature-based explanations might be effective in mitigating phishing attacks (Guidotti2019Survey; Gunning2019DARPAs). The clear advantage of such explanations is the reduction of susceptibility to phishing, but they are hand-made, expensive to maintain, and might not easily adjust to the changing circumstances of the threats (desolda2023explanations; buono2023warnings; greco2025enhancing). Other ways of generating explanations have also been considered in general, such as the use of counterfactuals (Guidotti2019Counterfactual) or the use of Large Language Models (LLMs), but they have not been investigated extensively in the specific field of cybersecurity (Xu2025LLMsCybersecurity).
@@
To bridge these gaps, we have carried out a large-scale, between-subjects user study ($N=750$). We investigated how well phishing warning dialogues with explanations can be generated with LLMs. We assessed the result of using a commercial closed-weight model (Claude 3.5 Sonnet) and an open-weight alternative (Llama 3.3 70B). Additionally, we considered two styles of explanation: feature-based and counterfactual explanations. A baseline with feature-based, manually written explanations has also been considered  (desolda2023explanations; greco2025enhancing). The effectiveness of the different experimental conditions has been evaluated by considering both measures of behaviour (e.g., click-through rate on phishing links) and user perceptions (e.g., clarity, trust, perceived risk), and by modelling the degree to which individual factors, including cognitive style and knowledge of cybersecurity issues, influence results.
@@
We found that explanations generated by LLMs can be as effective as those produced manually; specifically, explanations generated by Claude were better than those crafted manually. Regarding the type of explanation, counterfactual ones have an interesting power in preventing false positive cases, while feature-based ones have certain benefits in identifying true positive cases. Moreover, user-related factors such as gender, familiarity, and mental workload significantly influence the effectiveness of the warning.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on phishing defences, warning dialogue design, and explainable AI in cybersecurity. Section 3 presents the experimental design and the explanation generation process. This section also reports the formalisation of the four research questions this study aims to answer. Section 4 details the study procedure and methodology. Section 5 reports the quantitative and qualitative results addressing our four research questions. Section 6 discusses threats to validity. Section 7 synthesises the findings as a set of lessons learned. Section  8 concludes the paper by summarising the main results and discussing potential areas for future research.
@@ 2.1 Warning dialogues for Phishing Defence
The complete elimination of phishing through automated filtering methods such as blocklists (gupta2014emerging) and AI-based methods Basit2021Survey is currently unfeasible. No existing tools can detect phishing content, such as websites or emails, with perfect accuracy (elassal2020benchmarking; gholampour2023adversarial). Furthermore, attempts at reducing the number of false negatives (i.e., phishing emails that go undetected) tend to increase the number of false positives (i.e., legitimate emails misclassified as phishing) in classification tasks (saxena2018precision). This eventually leads to filtering out genuine, relevant emails, potentially disrupting user productivity.
@@
The solution currently adopted in most systems is not to automatically filter suspicious content classified as phishing with low confidence; instead, these emails are presented to users alongside a warning dialogue that alerts them to possible threats, leaving the final decision to the users (kumaraguru2010teaching). Despite warning dialogues, many users fail to grasp the significance or meaning of these messages, so phishing attacks remain highly effective (ibm2024security).
@@
One key problem of everyday warning dialogues is their passive behaviour. Popular email clients such as Gmail implement static toolbars above the email content. Warnings like these are largely ineffective, as users often overlook or ignore them (egelman2008warned; wu2006security). On the other hand, warnings that implement an active behaviour, i.e., interrupt user interactions and demand attention, have been proven to be significantly more effective (wogalter2002research; petelka2019put; buono2023warnings; egelman2008warned; greco2025enhancing). However, such warnings are currently implemented only in web browsers (desolda2019alerting), despite the benefits of active behaviour having also been demonstrated for warnings shown in email clients (greco2025enhancing; buono2023warnings).
@@
Another limitation of warnings is the phenomenon known as habituation. This behaviour occurs when users are repeatedly exposed to the same warning stimulus, even under varying risk conditions, leading to decreased attentional responses and increased likelihood of bypassing the warning (akhawe2013alice; kim2009habituation). To mitigate this problem, polymorphic warnings were demonstrated to be significantly more resistant to habituation than static warnings (anderson2015polymorphic). Such warnings are called polymorphic since they dynamically change their appearance or content to provide varied visual stimuli. Nonetheless, they are still not implemented in browsers or email clients despite their advantages (desolda2019alerting). Only a few studies have explored the use of polymorphic warnings paired with tailored explanations to enhance their effectiveness (desolda2023explanations; buono2023warnings), or visual nudges based on social salience to enhance effectiveness (nicholson2017can). desolda2023explanations found that including manually designed explanations of malicious features in warning dialogues in web browsers led to warnings that were more familiar and understandable than those in the most common browsers. buono2023warnings studied warnings with explanations in the context of email clients and found that active warnings with explanations based on phishing features largely outperformed contextual warnings without explanations. This study confirmed the importance of interrupting user interaction flow to focus entirely on the message, similar to active warnings shown in browsers. Additionally, the study demonstrated the importance of explaining to users why content is malicious, helping them make better-informed decisions.
@@
A further issue that compromises the effectiveness of the warnings is the absence of detailed explanations about the phishing risk (bravolillo2011improving). Ideally, warnings should communicate clearly why specific content is dangerous and outline the potential consequences of ignoring it. However, many current warnings provide vague messages that fail to identify the malicious elements within emails or websites. This places the burden of risk assessment on the users, who often lack cybersecurity expertise or additional information to make informed decisions (desolda2021human). This is particularly critical in lateral phishing scenarios, where users struggle to detect subtle changes in message content and over-rely on sender identity, which appears legitimate chitare2023may. By incorporating explanations into warnings, the users’ understanding of the threat can be improved, addressing knowledge gaps, and potentially increasing compliance and trust in the system (bravolillo2011improving; RN3). This aligns with findings suggesting that transparent communication about incidents is crucial for building trust and understanding between security practitioners and employees (chitare2025exploring).
@@
The use of explanatory warnings in web browsers has been explored in (desolda2023explanations). The authors identified seven website features that can be intuitively explained to users and produced two explanation messages for each. These messages were refined through an iterative process to optimise text readability and sentiment. Moreover, the explanation messages followed a well-defined structure grounded in warning theory (bauer2013warning), which led to the identification of a template for generating explanation messages: “Feature description + Hazard Explanation + Consequences of not complying with the warning”. The authors empirically validated the adoption of warnings with explanations among 150 users, comparing it with warnings without explanations. Results showed that the proposed warnings were clearer and more effective than those implemented in modern browsers, which lack explanations. greco2025enhancing explored a similar approach in the context of email clients, combining active warning dialogues with embedded explanations. The findings indicated that this combination of active behaviour and explanation significantly outperformed the state-of-the-art email client warnings proposed by petelka2019put.
@@
As security software engineers, we must understand the underlying decision-making algorithm to generate explanations. In the case of AI-based phishing filtering methods, explanations can be produced by employing eXplainable AI (XAI) techniques (Guidotti2019Survey; Gunning2019DARPAs). [^1] See (ViganoMagazzeni2020) for a discussion on the relationship between explainable cybersecurity and explainable AI. These allow one to explain the behaviour and decisions of AI models, which are often black-box, and thus not directly interpretable even by AI engineers. With XAI techniques, for example, the system can provide a rationale for the reasons that led to the automatic classification of an email as phishing. Explanations in warning dialogues can contain such a rationale to improve the transparency of the AI system. Let us explore these aspects in the following subsection.
@@ 2.2 Human-Centred Explainable AI for Phishing Attacks
Integrating XAI solutions into cybersecurity systems is crucial for maintaining user trust and confidence. This approach contrasts with other human-centric interventions, such as simulated phishing campaigns, which have been argued to potentially damage the trust relationship between employees and security teams Murdoch2017. This underscores the importance of user-centred experimental evaluations that balance security, usability, and resilience to adversarial attacks (Charmet2022XAI4Cybersecurity; Rjoub2023XAI4Cybersecurity; SARKER2024XAI4CS; Samed2025XAI4CSIntrusionDetectionSystems). In this regard, the broader Human-Centred XAI (HCXAI) literature offers valuable insights into multiple explanation types and their effects on users. The most prevalent explanations in AI-assisted decision systems are feature-based ones (Lai2023SurveyHumanAIDecisionMaking), which clarify why an AI system produced a specific decision by highlighting the most important features driving the decision. Feature-based explanations may enhance user understanding, awareness of uncertainty, trust calibration, and decision accuracy, although they can sometimes foster overreliance on AI (Zhang2020ConfidenceExplanationsAccuracyTrust; Wang2021ExampleBasedFeatureBasedAndOthers; Chen2023RelianceExampleBasedFeatureBased; Ma2023CorrectnessLikelihoodAIUsers; Fok24FeatureBased; Cau25MultiExp). Counterfactual explanations constitute another, less explored style: they provide not only why an AI system reached a specific decision, but also how it is possible for users to alter input features to achieve a potential desired outcome by presenting contrastive “what-if” statements. While they are mainly used to offer actionable recourse when negative decisions occur—e.g., loan denial—by indicating necessary feature changes for a positive result—e.g., increasing income for approval—(Koh24AlgorithmicRecourse; Verma24AlgorithmicRecourse; VanNostrand24CounterfactualActionableRecourse; Upadhyay25AlgorithmicRecourse; Cau2025HCXAILoan), they have also demonstrated improvements in helpfulness, decision alignment, and accuracy, sometimes matching or exceeding feature-based methods, in AI-assisted decision-making tasks (Scharowski2023FeatureImportanceCounterfactualsExplanations; Celar2023CounterfactualCausal; Cau2023LogicalReasoningStock; Teso2023InteractiveExplanations; Lee23Counterfactual; Gentile25Counterfactual).
@@
Regarding XAI in cybersecurity and phishing in particular, only a small number of proof-of-concept studies have assessed the effectiveness of explanations with real users (desolda2023explanations; greco2025enhancing). For example, desolda2023explanations found that adding a feature-based explanation to a phishing warning dialogue made the warning more understandable and familiar, which may deter users from visiting malicious sites. In a similar study, greco2025enhancing compared warnings with and without feature-based explanations, as well as the effect of displaying the warning before versus after opening an email. They demonstrated that adding explanations enhanced users’ comprehension of the phishing threat and, when the warning appeared after an email was opened, resulted in a measurable reduction in click-through rate (CTR) despite the additional effort required by participants. However, these investigations have focused exclusively on manually crafted feature‐based explanations, leaving the generation of feature‐based or alternative explanation types via large language models largely unexplored (SARKER2024XAI4CS).